OpenStudy (anonymous):

Solve the triangle. A=32 degrees, a=19, b=14

OpenStudy (jdoe0001):

what's there to "solve"?

OpenStudy (anonymous):

Sides B and C and angle c

OpenStudy (jdoe0001):

are you doing trig I assume? have you done the Law of Sines yet?

OpenStudy (anonymous):

That's what we are doing now

OpenStudy (jdoe0001):

ok, then

OpenStudy (jdoe0001):

$$ \cfrac{sin(B)}{14}=\cfrac{sin(32)}{19}\implies sin(B) =\cfrac{14sin(32)}{19} $$

OpenStudy (dumbcow):

 solve for angle B using law of sines \[\frac{\sin B}{14} = \frac{\sin 32}{19}\] from there you get angle C \[C = 180 - 32-B\] then solve for side c using law of sines again \[\frac{c}{\sin C} = \frac{19}{\sin 32}\]

OpenStudy (jdoe0001):

so, the rational above, as dumbcow showed, will give you a number, then arccosine bothe sides to get the angle B and the angle C will be 180-(A+B)

OpenStudy (jdoe0001):

as as dumbcow showed above, use the Law of Sines, once you have the angle C, to get the side "c" :)